# Scaling Multi-Agent Epistemic Planning through GNN-Derived Heuristics

### v2

@@ Abstract.
Multi-Agent Epistemic Planning, Graph-Neural Networks, Learning in Planning, Heuristics1. Introduction
Planning scenarios involving multiple interacting entities, referred to as multi-agent, have gained increasing importance due to their relevance in real-world applications, where groups of agents frequently need to interact. However, effectively addressing multi-agent settings poses one of the most interesting challenges in modern AI research: adequately modeling multi-agent interaction while maintaining tractability (Brafman and Domshlak, 2013). This is because such modeling requires accounting not only for the state of the world, but also for the dynamics of information exchange between agents. Such reasoning, which deals with formalizing belief relationships among multiple agents, is referred to as epistemic reasoning (Fagin et al., 1995).
Interest in Multi-agent Epistemic Planning (MEP)—which integrates epistemic reasoning with automated planning—has surged (Baral et al., 2017), and several epistemic planners have been proposed (Bolander and Andersen, 2011; Muise et al., 2015; Engesser et al., 2017; Kominis and Geffner, 2017; Huang et al., 2017; Fabiano et al., 2020; Bolander et al., 2021; Pham et al., 2023). To the best of our knowledge, only a few systems (Fabiano et al., 2020; Burigana et al., 2023a; Pham et al., 2023) are capable of reasoning over this setting without restrictions. Nonetheless, these systems are severely limited by high computational costs, often making solving impractical. This inefficiency stems mainly from two factors: (1) the intrinsic complexity of the underlying representations, which makes applying transitions and evaluating formulas within epistemic states (e-states) substantially harder than in classical planning; and (2) the lack of effective heuristics, which results in a blind, combinatorial search as plan length increases. While the aforementioned works in MEP largely address the first issue, few efforts tackle the latter. A notable exception is the -EFP planner (Fabiano et al., 2024), an extension of Le et al. (2018), which integrates heuristics guidance to improve scalability. Our work builds upon this direction, sharing the core objective of designing effective heuristics extraction methods. We argue this focus is essential, as informed search is what enables scalability in planning systems—from classical heuristics planning (Bonet and Geffner, 2001; Helmert, 2006) to Monte Carlo Tree Search (MCTS) in reinforcement learning (RL) (Bouzy and Chaslot, 2006).
@@
The remainder of this paper is structured as follows. In Section 2, we provide background on MEP and GNNs. Section 3 presents our main theoretical contribution. In particular, Section 3.1 presents the design of the embedding and the dataset generation while Section 3.4 illustrate the training of the GNN-based regressor. Section 4 reports experimental results that evaluate the performance and scalability of our approach. We discuss limitations and related work in Sections 5 and 6, and conclude in Section 7.
@@ 2. Background
2.1. Epistemic Reasoning
2.1.1. Dynamic Epistemic Logic.
@@ 2.1. Epistemic Reasoning
Let us denote as a set of agents such that with , and as a set of propositional variables, referred to as fluents literals, or simply fluents. Each world is described by a subset of elements from intuitively, those deemed True. Furthermore, in epistemic logic, each agent is associated to an epistemic modal operator , signifying the belief222We use the terms knowledge and belief interchangeably, as their distinction is beyond this work’s scope. See Fagin et al. (1995) for a full discussion. of the agent. Additionally, the epistemic group operator is introduced. Essentially, this operator represents the common knowledge of a group of agents .
To be more precise, as in Baral et al. (2022), we have that a fluent formula is a propositional formula built using fluents in as propositional variables and the propositional operators . On the other hand, a belief formula is either (i) a fluent formula; (ii) if is a belief formula and , then is a belief formula; (iii) if and are belief formulae, then and are belief formulae, where ; or (iv) if is a belief formula and then is a belief formula. denotes the language of the belief formulae over the set .
The classical way of providing semantics for epistemic logic is in terms of pointed Kripke structures (Kripke, 1963).
@@
Intuitively, to verify whether a belief formula holds, we need to apply reachability within the Kripke model representing the e-state. By exploring the set of reachable worlds obtained by applying epistemic operators, we determine which configurations of fluents an agent (or group of agents) considers possible. Inconsistencies among these reachable worlds are used to model ignorance. The formal semantics over pointed Kripke structures is provided in Fabiano et al. (2020); Baral et al. (2022) and omitted here as it is not integral to understanding the contribution of this paper.
2.1.2. Multi-Agent Epistemic Planning.
In this section, we will introduce the fundamental concepts relevant to MEP, while addressing interested readers to Fagin et al. (1995); Bolander and Andersen (2011) for a more exhaustive introduction.
Let us begin by defining the notion of a multi-agent epistemic planning problem in Definition 2.1.2. Intuitively, an epistemic planning problem encompasses all the necessary information to frame a planning problem within a multi-agent scenario.
@@
To the best of our knowledge, the most accepted formalization of a comprehensive action language for multi-agent epistemic planning is (Baral et al., 2022). Note that other languages capable of reasoning about DEL also exist (Muise et al., 2015; Bolander et al., 2021), but they limit their expressiveness in favor of efficiency and are therefore not considered here.
@@
In their work Baral et al. (2022) delineate three distinct types of actions within the context of multi-agent domains: (i) world-alteringactions (or ontic actions): employed to modify specific properties, or fluents, within the world—denoted through the statement “ causes 333 can be either a fluent or its negation”; (ii) sensingactions: used by an agent to refine her beliefs about the world—denoted through the statement “ determines ”; and (iii) announcementactions: utilized by an agent to influence the beliefs held by other agents—denoted through the statement “ announces ”. For brevity, we will not provide further details of here. Interested readers can find a comprehensive description in Baral et al. (2022).
2.2. Graph Neural Networks
@@
Formally, let be a graph with nodes, adjacency matrix (augmented with self‑loops as ), and node‐feature matrix . A Graph Convolutional Network (GCN) Kipf and Welling (2017) layer then computes
where , are learnable weight matrices, and is an activation function. The purpose is to derive informative vector representations (node embeddings) serving as concise, low-dimensional summaries capturing a node’s data, its relational context, and its role within the broader graph topology.
@@ 3. Learning MEP Heuristics with GNN
As discussed, planning in MEP is an extremely resource-intensive task. To mitigate this computational burden, we explore alternatives to Breadth-First Search (BFS), such as Best-First Search, which we abbreviate as HFS (for Heuristics-First Search) to distinguish it from BFS. HFS is a widely used strategy that prioritizes expanding states with the most favorable heuristics score. Both BFS and HFS are standard approaches described in Russell and Norvig (2020, Ch. 11).
BFS is an uninformed search method, exploring the state space uniformly. In contrast, HFS aims to guide the search more efficiently by prioritizing e-states that are likely closer to the goal. The key challenge lies in defining an effective evaluation function—also known as a heuristics—capable of accurately ranking epistemic states. Heuristics are a central topic in the planning literature and have been extensively studied and proven effective (Keyder and Geffner, 2008; Russell and Norvig, 2020). For this reason, a key contribution of this work is the formalization of heuristics tailored for MEP through the employment of GNNs.
@@ 3.1. Training Data Generation
The first key challenge in employing ML techniques in MEP is determining what kind of information can be meaningfully extrapolated from data. Several approaches have been proposed in automated planning, such as training models end-to-end using full problem descriptions and their associated goals (Pallagani et al., 2023; Huang et al., 2024). However, these methods often suffer from low accuracy and require a large number of training instances to generalize effectively (Kambhampati et al., 2024).
To address these limitations, we adopt a learning approach at the level of individual e-state, treating each state as an independent training signal rather than relying on complete trajectories. This offers two main advantages. First, although individual predictions may occasionally be inaccurate, their impact on the overall search is limited—as long as the trend of the heuristics is informative, the search remains effective. Second, this approach significantly reduces the amount of data required. Instead of needing complete problems as training instances, we can treat each e-state encountered during exploration as a distinct data point. This allows a single run of the planner to generate tens of thousands of e-states, yielding a large set of training examples.
The main objective of this work is to employ GNNs to extract information from this data to guide the search as if we had access to the perfect heuristics—the function that assigns to each state its distance to the nearest goal. Coupled with the HFS search algorithm, this heuristics permits navigating the search space optimally, returning one of the shortest plans while expanding the minimal number of nodes.
@@
The next challenge we addressed was designing a suitable data representation for integration within a GNN. Each e-state is formalized as a pointed Kripke structure (Definition 2.1.1), namely
@@
We explored three strategies, each balancing information preservation and efficiency differently.
@@
Each world is assigned a symbolic integer identifier through a mapping Intuitively, this assigns consecutive integers to newly observed worlds. However, since this assignment depends on the order in which worlds are generated, is not invariant across runs, leading to inconsistencies between datasets. This representation considerably reduces data generation, training, and inference time but incurs the highest degree of information loss, as similarities between worlds are not preserved.
@@
A hash function is applied to each world: ; where denotes the repetition index distinguishing identical copies of the same world within an e-state. This ensures consistent world identifiers across runs, but relative distances between worlds are lost, as, for example, worlds differing only by a single fluent value may receive completely unrelated hash values. This encoding represents a middle ground in terms of the trade-off between information loss and the computational power required.
@@
Each world is represented as a purely binary vector: where is the binary representation of the repetition index , and if fluent holds in , and otherwise (or when ). Note that, by construction, is assumed to be greater than . Thus, encodes both the repetition number of the world within the e-state and its truth assignment over as a single binary vector. This embedding preserves all logical information about the e-estate, but it does so at the cost of larger encodings and slower generation, training, and inference.
We compare these three representations in Section 4.1, showing that hash-based encoding achieves the best performance.
@@ 3.2.1. Goal Encoding within the State Embedding
While the encoded Kripke model—represented through the dot language (Ellson et al., 2004)—proved effective, we encountered a second challenge. Training solely on states led the regressor to learn absolute measures of distance that did not depend on the underlying goal, generalizing only to problems with identical goal conditions. This would limit the applicability of the heuristics to instances sharing the same goal.
@@
In the resulting formulation, two types of nodes are encoded for each e-state embedding: (i) state nodes, corresponding to worlds in the epistemic state; (ii) goal nodes, corresponding to symbols describing the goal conditions. This is exemplified in Figure 1, where the left portion of the expanded hashed-based e-state (in teal) shows the goal encoding, and the right portion (in magenta) depicts the e-state itself. The two are connected via a shared graph structure using nodes and edges that preserve constants and special identifiers throughout the process.
In the hashed- and in the mapping-based representations, the state component remains unchanged, while goal nodes are associated with unique integer identifiers (assigned to the first occurrences of each goal symbol). In the bitmask representation, both node types share the same binary vector structure: where is reflects the bitmask embedding presented above and is the binary segment reserved for goal encoding. During experimental evaluation, we fixed the total size of to 64 bits, with , , and .
@@ 3.3. Building the Dataset
Although conceptually simple, this process faces a combinatorial explosion: even with just 10 actions and a depth of 25—an overestimation that captures the typical plan length in standard MEP benchmarks—the search space can reach nodes, making exhaustive exploration infeasible. To address this, we draw on ideas from local search (Gerevini et al., 2003), sampling different subregions of the space to try and maximize coverage. Our DFS includes a branch discard probability (adaptive to depth and node count), randomized action ordering to avoid prefix bias, a cap on the number of explored nodes, and duplicate e-state checks. These mechanisms allow us to explore the search space in a sufficiently diverse yet tractable manner, enabling the construction of informative datasets within minutes per instance.
While data generation and training are handled offline in this work, the structure of our learning pipeline naturally supports an online setup. By slightly extending the search process, the planner could incrementally collect training pairs and update the GNN once a sufficient number of samples is accumulated. Thanks to deep’s multithreading support—introduced to emulate the portfolio behavior of -EFP (Fabiano et al., 2024)—this online learning loop could run in parallel, allowing the planner to adapt and improve over time, similar to the cognitive architecture presented in Fabiano et al. (2025). We leave the exploration of this online learning paradigm to future work.
@@ 3.4. Training Neural Distance Estimator
Data Preparation.
Our objective is to train an estimator that, given an e-state predicts the distance to a goal state.
To effectively train the estimator, three design choices were made:
@@
Linearly normalize the distance target. To stabilize training, we linearly normalize the true distance from to , with . Let
Then the normalized target is
(1) In inference, we recover the original scale through
@@
Training Process.
Our neural regressor is implemented in PyTorch Imambi et al. (2021), and its graph encoder is built using PyTorch Geometric Bielak and Kajdanowicz (2022). For each epistemic state (e-state), represented as a graph with nodes and edges , we construct a corresponding data object that encodes the graph structure. This object serves as input to the GNN, which processes it to produce a latent embedding. Specifically, this object includes the following: (i) Node identifiers:, represented as bit-vectors, in which each node is expressed as a binary vector , where is the bit-length (here ). The resulting node feature matrix is ; (ii) Edge indices:, encoded as an index tensor , whose columns list each source–target pair; (iii) Edge attributes:, stacked into an attribute tensor and normalized to . The estimator is trained by minimizing the Mean Squared Error () between its predicted scalar and the true normalized distance : using the AdamW optimizer with the parameters listed in Appendix B.
@@
Bitwise embedding of node identifiers and edge attributes: Each node bit-vector is projected through a two-layer MLP,
producing dense embeddings of dimension . Similarly, each edge attribute is processed through
yielding -dimensional embeddings. This bitwise encoding provides a compact and lossless representation of node identifiers while supporting continuous optimization and stable gradients.
Relational message passing with GINEConv: Two successive GINEConv (Hu et al., 2020) layers, each followed by a ReLU activation, allow each node to aggregate information from its neighbors while explicitly incorporating edge indices and attributes. Formally, the GINE update for node in layer is
[…]
Graph-level summarization: The node embeddings from the final GINE layer are aggregated into a fixed-size representation via mean pooling:
This operation is permutation-invariant and scale-independent, ensuring stable representations across graphs of different sizes. In contrast, sum pooling scales with , potentially biasing larger graphs, while attention pooling introduces learnable weights but increases variance and computational cost. Mean pooling therefore offers the best generalization–variance trade-off for distance regression.
Deep residual regression head with bounded output: The pooled graph embedding is processed by a deep residual regressor:
where denotes a sequence of ResidualBlocks (Linear BatchNorm ReLU Dropout Linear skip). The sigmoid bounds the output in , and is finally clamped to to prevent numerical instabilities near the boundaries.
Overall, the combination of bitwise node embeddings, GINEConv message passing, and mean pooling yields a compact and expressive model that generalizes across variable-sized epistemic graphs while remaining numerically stable during training. The full forward-pass procedure is detailed in Appendix E.

### v3

@@ Abstract.
Multi-Agent Epistemic Planning; Graph-Neural Networks; Learning in Planning; Heuristicsdoi:
QMRY96611. Introduction
Planning scenarios involving multiple interacting entities, referred to as multi-agent, have gained increasing importance due to their relevance in real-world applications, where groups of agents frequently need to interact. However, effectively addressing multi-agent settings poses one of the most interesting challenges in modern AI research: adequately modeling multi-agent interaction while maintaining tractability (BRAFMAN201352). This is because such modeling requires accounting not only for the state of the world, but also for the dynamics of information exchange between agents. Such reasoning, which deals with formalizing belief relationships among multiple agents, is referred to as epistemic reasoning (fagin1994reasoning).
Interest in Multi-agent Epistemic Planning (MEP)—which integrates epistemic reasoning with automated planning—has surged (BBMvD17), and several epistemic planners have been proposed (BolanderA11; muise2015planning; EngesserBMN17; KominisG17; huang2017general; icaps20; KR2021-12; Pham_Son_Pontelli_2023). To the best of our knowledge, only a few systems (icaps20; DBLP:conf/jelia/BuriganaFM23; Pham_Son_Pontelli_2023) are capable of reasoning over this setting without restrictions. Nonetheless, these systems are severely limited by high computational costs, often making solving impractical. This inefficiency stems mainly from two factors: (1) the intrinsic complexity of the underlying representations, which makes applying transitions and evaluating formulas within epistemic states (e-states) substantially harder than in classical planning; and (2) the lack of effective heuristics, which results in a blind, combinatorial search as plan length increases. While the aforementioned works in MEP largely address the first issue, few efforts tackle the latter. A notable exception is the -EFP planner (DBLP:conf/prima/FabianoPSP24), an extension of le2018efp, which integrates heuristics guidance to improve scalability. Our work builds upon this direction, sharing the core objective of designing effective heuristics extraction methods. We argue this focus is essential, as informed search is what enables scalability in planning systems—from classical heuristics planning (BONET20015; 10.5555/1622559.1622565) to Monte Carlo Tree Search (MCTS) in reinforcement learning (RL) (bouzy2006monte).
@@
The remainder of this paper is structured as follows. In Section 2, we provide background on MEP and GNNs. Section 3 presents our main theoretical contribution. In particular, Section 3.1 presents the design of the embedding and the dataset generation while Section 3.4 illustrates the training of the GNN-based regressor. Section 4 reports experimental results that evaluate the performance and scalability of our approach. We discuss limitations and related work in Sections 5 and 6, and conclude in Section 7.
@@ 2. Background
2.1. Dynamic Epistemic Logic
@@
Let us denote as a set of agents such that with , and as a set of propositional variables, referred to as fluents literals, or simply fluents. Each world is described by a subset of elements from intuitively, those deemed True. Furthermore, in epistemic logic, each agent is associated to an epistemic modal operator , signifying the belief222We use the terms knowledge and belief interchangeably, as their distinction is beyond this work’s scope. See fagin1994reasoning for a full discussion. of the agent. Additionally, the epistemic group operator is introduced. Essentially, this operator represents the common knowledge of a group of agents .
To be more precise, as in baral2021action, we have that a fluent formula is a propositional formula built using fluents in as propositional variables and the propositional operators . On the other hand, a belief formula is either
a fluent formula;
if is a belief formula and , then is a belief formula;
if and are belief formulae, then and are belief formulae, where ; or
if is a belief formula and then is a belief formula.
denotes the language of the belief formulae over the set .
The classical way of providing semantics for epistemic logic is in terms of pointed Kripke structures (Kripke1963-KRISCO).
@@
Intuitively, to verify whether a belief formula holds, we need to apply reachability within the Kripke model representing the e-state. By exploring the set of reachable worlds obtained by applying epistemic operators, we determine which configurations of fluents an agent (or group of agents) considers possible. Inconsistencies among these reachable worlds are used to model ignorance. The formal semantics over pointed Kripke structures is provided in icaps20; baral2021action and omitted here as it is not integral to understanding the contribution of this paper.
2.2. Multi-Agent Epistemic Planning
We are now ready to introduce the fundamental concepts of MEP, while addressing interested readers to fagin1994reasoning; BolanderA11 for a more exhaustive introduction.
Let us begin by providing the notion of a multi-agent epistemic planning problem in Definition 2.2. Intuitively, an epistemic planning problem encompasses all the necessary information to frame a planning problem within a multi-agent scenario.
@@ 2.2. Multi-Agent Epistemic Planning
To the best of our knowledge, the most accepted formalization of a comprehensive action language for multi-agent epistemic planning is (baral2021action). Note that other languages capable of reasoning about DEL also exist (muise2015planning; KR2021-12), but they limit their expressiveness in favor of efficiency and are therefore not considered here.
@@
In their work baral2021action delineate three distinct types of actions within the context of multi-agent domains: (i) world-alteringactions (or ontic actions): employed to modify specific properties, or fluents, within the world—denoted through the statement “ causes 333 can be either a fluent or its negation”; (ii) sensingactions: used by an agent to refine her beliefs about the world—denoted through the statement “ determines ”; and (iii) announcementactions: utilized by an agent to influence the beliefs held by other agents—denoted through the statement “ announces ”. For brevity, we will not provide further details of here. Interested readers can find a comprehensive description in baral2021action.
2.3. Graph Neural Networks (GNNs)
@@
In our neural estimator, the GNN forms the first stage of the pipeline: it processes the graph representation of an epistemic state and produces a compact latent embedding that serves as the basis for the heuristic estimate, capturing the structural and logical characteristics of the state. Here, message passing is implemented via GINEConv (DBLP:conf/iclr/HuLGZLPL20): given a node and its neighborhood , the update at layer is
[…]
Reviews of architectures and applications are proposed in wu2020comprehensive and zhou2020graph, respectively.
@@ 3. Learning MEP Heuristics with GNN
As discussed, planning in MEP is an extremely resource-intensive task. To mitigate this computational burden, we explore alternatives to Breadth-First Search (BFS), such as Best-First Search, which we abbreviate as HFS (for Heuristics-First Search) to distinguish it from BFS. HFS is a widely used strategy that prioritizes expanding states with the most favorable heuristics score. Both BFS and HFS are standard approaches described in modernApproach.
BFS is an uninformed search method, exploring the state space uniformly. In contrast, HFS aims to guide the search more efficiently by prioritizing e-states that are likely closer to the goal. The key challenge lies in defining an effective evaluation function—also known as a heuristics—capable of accurately ranking epistemic states. Heuristics are a central topic in the planning literature and have been extensively studied and proven effective (Keyder:2008:HPA:1567281.1567409; modernApproach). For this reason, a key contribution of this work is the formalization of heuristics tailored for MEP through the employment of GNNs.
@@ 3.1. Training Data Generation
The first key challenge in employing ML techniques in MEP is determining what kind of information can be meaningfully extrapolated from data. Several approaches have been proposed in automated planning, such as training models end-to-end using full problem descriptions and their associated goals (ijcai2023p839; huang2024planningdarkllmsymbolicplanning). However, these methods often suffer from low accuracy and require a large number of training instances to generalize effectively (kambhampati2024llmscantplanhelp).
To address these limitations, we adopt a learning approach in which every state represents an independent training signal, rather than relying on complete trajectories. This offers two main advantages. First, although individual predictions may occasionally be inaccurate, their impact on the overall search is limited—as long as the trend of the heuristics is informative, the search remains effective. Second, this approach significantly reduces the amount of data required. Instead of needing complete problems as training instances, we can treat each e-state encountered during exploration as a distinct data point. This allows the generation of tens of thousands of e-states, yielding a large set of training examples in a single run of the planner.
The main objective of this work is to use GNNs to extract information from the data to guide the search. In particular, we aim at approximating the perfect heuristics that assign to each state its distance to the nearest goal.
@@
The next challenge we addressed was designing a suitable data representation for integration within a GNN. Each e-state is formalized as a pointed Kripke structure (Definition 2.1), namely
@@
We explored three strategies, each offering a different trade-off between information preservation and efficiency. In Section 4.1, we compare these representations, illustrated below, and show that hash-based encoding achieves the best performance.
@@
Each world is assigned a symbolic integer identifier through a mapping Intuitively, this assigns consecutive integers to newly observed worlds. However, since this assignment depends on the order in which worlds are generated, making not invariant across runs, leading to inconsistencies between datasets. This representation has the lowest computational burden for data generation, training, and inference, but it incurs the greatest information loss, as similarities between worlds are not preserved.
@@
A hash function is applied to each world: ; where denotes the repetition index distinguishing identical copies of the same world within an e-state. This ensures consistent world identifiers across runs, but relative distances between world evaluations are lost, as, for example, evaluations differing only by a single fluent value may receive completely unrelated hash values. We use the hash_range function from the Boost libraries (10.5555/2049814) due to its high performance and existing use in deep for e-state storage, which minimizes computational overhead. Alternative hashing approaches may offer further improvements and are left for future work.
This encoding represents a middle ground in terms of the trade-off between information loss and the computational power required.
@@
Each world is represented as a purely binary vector: where is the binary representation of the repetition index , and if fluent holds in , and otherwise (or when ). Note that, by construction, is assumed to be greater than . Thus, encodes both the repetition number of the world within the e-state and its truth assignment over as a single binary vector. This embedding preserves most of the logical information about the e-estate, but it comes at the cost of a larger encoding, which requires more data for the model to converge appropriately and results in slower generation, training, and inference.
@@ 3.2.1. Goal Encoding within the State Embedding
While the encoded Kripke model—represented through the dot language (Ellson2004)—proved effective, we encountered a second challenge. Training solely on states led the regressor to learn absolute measures of distance. This measure is independent of the underlying goal and therefore generalizes only to problems with identical goal conditions, which limits the applicability of the heuristics.
@@
In the resulting formulation, two types of nodes are encoded for each e-state embedding: (i) state nodes, corresponding to worlds in the epistemic state; (ii) goal nodes, corresponding to symbols describing the goal conditions. This is exemplified in Figure 1, where the left portion of the expanded hashed-based e-state (in teal) shows the goal encoding, and the right portion (in magenta) depicts the e-state itself. The two are connected via a shared graph structure using nodes and edges that use constants and special identifiers throughout the process.
In the hashed- and in the mapping-based representations, the state component remains unchanged, while goal nodes are associated with unique integer identifiers (assigned to the first occurrences of each goal symbol). In the bitmask representation, both node types share the same binary vector structure: where reflects the bitmask embedding presented above and is the binary segment reserved for goal encoding. During the experimental evaluation, we fixed the total size of to 64 bits, with , , and .
@@ 3.3. Building the Dataset
Although conceptually simple, this process suffers from combinatorial explosion: with only 10 actions and depth 25—an overestimate of typical plan length in standard MEP benchmarks—the search space can reach nodes, making exhaustive exploration infeasible. To mitigate this, we draw on ideas from local search (Gerevini_2003), sampling subregions to maximize coverage. Our DFS incorporates probabilistic branch pruning (adaptive to depth and node count), randomized action ordering to avoid prefix bias, a node expansion cap, and duplicate e-state checks. These mechanisms enable diverse yet tractable exploration, allowing informative datasets to be built within minutes per instance.
While data generation and training are handled offline in this work, the structure of our learning pipeline naturally supports an online setup. By slightly extending the search process, the planner could incrementally collect training pairs and update the GNN once a sufficient number of samples is accumulated. Thanks to deep’s multithreading support—introduced to emulate the portfolio behavior of -EFP (DBLP:conf/prima/FabianoPSP24)—this online learning loop could run in parallel, allowing the planner to adapt and improve over time, similar to the cognitive architecture presented in 10.1145/3715709. We leave the exploration of this online learning paradigm to future work.
@@ 3.4. Training Neural Distance Estimator
Our objective is to train an estimator that, given an e-state, predicts the distance to a goal state. To this end, three design choices were made:
@@
Linearly normalize the distance target. To stabilize training, we linearly normalize the true distance from to , with . Let then the normalized target is In inference, we recover the original scale through
@@
Our neural regressor is implemented in PyTorch imambi2021pytorch, and its graph encoder is built using PyTorch Geometric bielak2022pytorch. For each epistemic state (e-state), represented as a graph with nodes and edges , we construct a corresponding data object that encodes the graph structure. This object serves as input to the GNN, which processes it to produce a latent embedding. Specifically, this object includes the following:
Node identifiers: , represented as bit-vectors, in which each node is expressed as a binary vector , where is the bit-length (here ). The resulting node feature matrix is ;
Edge indices: , encoded as an index tensor , whose columns list each source–target pair;
Edge attributes: , stacked into an attribute tensor and normalized to .
The estimator is trained by minimizing the Mean Squared Error () between its predicted scalar and the true normalized distance
using the AdamW optimizer with the parameters listed in Appendix B, and where is the mini-batch size (number of samples per gradient update).
@@
Bitwise embedding of node identifiers and edge attributes: Each node bit-vector is projected through a two-layer MLP, producing dense embeddings of dimension . Similarly, each edge attribute is processed through yielding -dimensional embeddings. This bitwise encoding provides a compact and lossless representation of node identifiers while supporting continuous optimization and stable gradients.
Relational message passing with GINEConv: Two successive GINEConv (DBLP:conf/iclr/HuLGZLPL20) layers, each followed by a ReLU activation, allow each node to aggregate information from its neighbors while explicitly incorporating edge indices and attributes.
Graph-level summarization: The node embeddings from the final GINE layer are aggregated into a fixed-size representation via mean pooling: This operation is permutation-invariant and scale-independent, ensuring stable representations across graphs of different sizes. In contrast, sum pooling scales with , potentially biasing larger graphs, while attention pooling introduces learnable weights but increases variance and computational cost. Mean pooling therefore offers the best generalization–variance trade-off for distance regression.
Deep residual regression head with bounded output: The pooled graph embedding is processed by a deep residual regressor: where denotes a sequence of ResidualBlocks (Linear BatchNorm ReLU Dropout Linear skip). The sigmoid bounds the output in , and is finally clamped to to prevent numerical instabilities near the boundaries.
Overall, the combination of bitwise node embeddings, GINEConv message passing, and mean pooling yields a compact and expressive model that generalizes across variable-sized epistemic graphs while remaining numerically stable during training. The full forward-pass is detailed in Appendix E.